### v1

@@ 7.3.3 Signal Computations
In order to facilitate computations over signals, the SPARQL algebra was further extended with expressions that represent the and operators from Section 6. The proposed extension contains two versions of existing SPARQL functions and operators: one defined for RDF terms and a lifted one defined for RDF signals. The regular SPARQL functions and operators are lifted via point-wise application, as shown in Section 4. The evaluation of the lifted version is defined via . Query engines must automatically use the lifted SPARQL functions and operators if any of their arguments refer to a signal. As the only way to introduce signals is the SIGNALS clause, query engines can infer the necessary information by looking at the signal declarations and tracking the usage of the variables.
Whenever a lifted SPARQL operator or function is used, a query engine must insert expressions for non-signal arguments. This ensures that, for a given function or operator, all arguments are either RDF terms or RDF signals. The operators are inserted according to the regular SPARQL behavior.
@@ 7.3.4 An Example of Evaluating an Expression
Assume that a query contains the expression ?ap * ?sign where ?ap is a signal that refers to the ev:ActivePower of a ?device and ?sign refers to an RDF term that indicates whether the AP must be inverted (PV systems). Furthermore, assume that the result of the expression is bound to the ?real_ap variable. As one subexpression of ?ap * ?sign is a signal, the lifted version of * is used. Then, since ?sign is a regular RDF term, the ?sign variable is wrapped in a expression to ensure that all arguments are signals. Listing 5 depicts the nested operators for evaluating the expression.
@@ 7.3.5 Query Forms
SigSPARQL queries that have a SIGNALS clause enable continuous queries. The result of a continuous SELECT query is a temporal result set equal to the TSS of the outermost algebra operator of the query. This result set also includes bindings to the same signals. As the signals model temporal values, users can evaluate the signals in the result set at different time points analogously to evaluating a TSSs. The result of a continuous CONSTRUCT query is a continuously growing RDF graph, as a single solution can cause multiple trigger events. The result of a CONSTRUCT query is the set union of all triples by substituting the template with the solution variables in the trigger events.
In regular SPARQL, blank nodes in the template are scoped to every solution [21]. In continuous CONSTRUCT queries, blank nodes in the template are scoped to each trigger event. Therefore, blank nodes are scoped to each template substitution, similarly as in regular SPARQL with a solution-based substitution. This allows users, for example, to use blank nodes to describe operating envelope violations. As each violation has a distinct trigger event, the resulting blank nodes will be distinct due to the scoping rules.
@@ 8 Feasibility Evaluation
We have implemented the proposed SPARQL extension in a prototype333https://doi.org/10.5281/zenodo.15260651 based on Oxigraph [30] to demonstrate its technical feasibility. The implementation uses discrete time and models the data between observations using the “last observation” strategy. The evaluation of continuous queries is done in two steps.
@@
Then, this result can be registered in a continuous query engine that handles the processing of time series data by continuously updating the signals in the result set with arriving values. After a query has been registered, the temporal result set for SELECT queries can be evaluated at different time points, while the graph result of CONSTRUCT queries is automatically updated for every detected trigger event. The implementation contains a test suite with multiple queries in the context of the motivating example.
@@ 9 Proof of Concept
The query then obtains the ev:ActivePower signal for each ?device and binds the resulting signal to the ?ap variable (line 12). Then, all solutions (including the active power signals) are grouped according to the garage that contains them (line 20). After grouping, the signal that describes the ev:Envelope of the ?garage is bound to the ?env variable (line 12). The condition for firing trigger events is determined by the WHEN clause (lines 7-10), while the CONSTRUCT clause (lines 1-6) defines the triple template. When this query is continuously evaluated, the result is a continuously growing RDF graph that describes all envelope violations. Note that due to the scoping rules for blank nodes in the CONSTRUCT clause, each violation (trigger event) becomes a distinct node in the graph. In combination with Section 8, we have provided an answer to RQ2 by demonstrating that it is feasible to implement the proposed SPARQL extension and that it can address a typical CPS monitoring use case.
@@ 10 Discussion
The ability to use knowledge graphs to define continuous queries via signals has potential in monitoring CPS. This is because often two CPSs are slightly different from each other, even though they are in the same domain. For example, in the motivating example, an operator may own hundreds of garages. Although they will adhere to the same rules, many garages will be different (e.g., number of chargers). The proposed approach allows operators to formulate a single query that can be applied to all of their charging garages.
@@ 11 Conclusion & Future Work
In this work, we discuss why signals are an appropriate concept for formulating computations based on time series data in knowledge graph query languages and, in particular, SPARQL. We have identified three core operators that facilitate the integration between knowledge graphs and signals and propose SigSPARQL, a SPARQL extension that incorporates the identified operators and adapts the semantics of SELECT and CONSTRUCT queries to support continuously running queries. We demonstrated the technical feasibility of our approach by implementing a prototype and applying it to a monitoring use case within a CPS that is part of an ongoing research project.
Future work includes extending the proposed semantics to temporal knowledge graphs and extending the language to support multiple event detection methods, such as efforts from the runtime verification and CEP communities. Furthermore, the development of an optimized prototype that adheres to the proposed semantics would benefit further applied work and can be used for performance experiments.

### v2

@@ 7.3.3 Signal Computations
In order to facilitate computations over signals, the SPARQL algebra was further extended with expressions that represent the and operators from Section 6. The proposed extension contains two versions of existing SPARQL functions and operators: one defined for RDF terms and a lifted one defined for RDF signals. The regular SPARQL functions and operators are lifted via point-wise application, as shown in Section 4. The evaluation of the lifted version is defined via . Query engines must automatically use the lifted SPARQL functions and operators if any of their arguments refer to a signal. As the only way to introduce signals is the SIGNALS clause, query engines can obtain the necessary information by looking at the signal declarations and tracking the usage of the variables.
Whenever a lifted SPARQL operator or function is used, all of its non-signal arguments must be lifted by using a expression. This ensures that, for a given function or operator, all arguments are either RDF terms or RDF signals. The graph patterns are inserted according to the regular SPARQL behavior.
@@ 7.3.4 An Example of Evaluating an Expression
Assume that a query contains the expression ?ap * ?sign, where ?ap is a signal referring to the ev:ActivePower of a ?device and ?sign refers to an RDF term indicating whether AP must be inverted (for PV systems). Furthermore, assume that the result of the expression is bound to the ?real_ap variable. As one subexpression of ?ap * ?sign is a signal, the lifted version of * is used. Then, since ?sign is a regular RDF term, the ?sign variable is wrapped in a expression to ensure that all arguments are signals. Listing 5 depicts the nested graph patterns for the expression.
@@ 7.3.5 Query Forms
SigSPARQL queries that have a SIGNALS clause enable continuous queries. The result of a continuous SELECT query is a temporal result set equal to the TSS of the outermost graph pattern of the query. This result set also includes bindings to the same signals. As the signals model temporal values, users can evaluate the signals in the result set at different time points analogously to evaluating a TSSs. The result of a CONSTRUCT query is the set union of all triples by substituting the template with the solution variables in the trigger events. In other words, the result is a continuously growing RDF graph, as a single solution can cause multiple trigger events.
In regular SPARQL, blank nodes in the template are scoped to every solution [21]. In continuous CONSTRUCT queries, blank nodes in the template are scoped to each trigger event. Therefore, blank nodes are scoped to each template substitution, similarly as in regular SPARQL. This allows users, for example, to use blank nodes to describe operating envelope violations. As each violation has a distinct trigger event, the resulting blank nodes will be different due to the scoping rules. The query in Section 9 uses the adapted scoping rules to generate a distinct blank node for each violation.
@@ 8 Feasibility Evaluation
We have implemented the proposed SPARQL extension in a prototype222https://doi.org/10.5281/zenodo.15260651 based on Oxigraph [29] to demonstrate its technical feasibility. The implementation uses discrete time and models the data between observations using the “last observation” strategy. The evaluation of continuous queries is done in two steps.
@@
Then, this result can be registered in a continuous query engine. The continuous query engine can maintain the intermediate result of the query (i.e., the result of evaluating regular SPARQL constructs), including the expressions mentioned above. The continuous query engine then registers for updates on the underlying signals and processes incoming updates. The prototype facilitates this by inserting newly arriving values into a time series that holds the underlying data for the signal. Based on this time series, the signal models the values between the data points. These updates also trigger the evaluation of the operators in the query.
After a query has been registered, the temporal result set for SELECT queries can be evaluated at different time points, while the graph result of CONSTRUCT queries is automatically updated for every detected trigger event. The implementation contains a test suite with multiple queries in the context of the motivating example.
@@ 9 Proof of Concept
The query then obtains the ev:ActivePower signal for each ?device and binds the resulting signal to the ?ap variable (line 12). Then, all solutions (including the active power signals) are grouped according to the garage that contains them (line 20). After grouping, the signal that describes the ev:Envelope of the ?garage is bound to the ?env variable (line 12). The condition for firing trigger events is determined by the WHEN clause (lines 7-10), while the CONSTRUCT clause (lines 1-6) defines the triple template. The time point of each trigger event is bound to ?violation_time (line 9). This information can be used as the time stamp of the violation. While this query is running, the result is a continuously growing RDF graph that describes all envelope violations. Note that due to the scoping rules for blank nodes in the CONSTRUCT clause, each violation (trigger event) becomes a distinct node in the graph. In combination with Section 8, we have provided an answer to RQ2 by demonstrating that it is feasible to implement the proposed SPARQL extension and that it can address a typical CPS monitoring use case.
@@ 10 Discussion
The ability to use knowledge graphs to define continuous queries via signals has potential in monitoring CPS. This is because often two CPSs are slightly different from each other, even though they are in the same domain. For example, in the motivating example, an operator may own hundreds of garages. Although they will adhere to the same rules, many garages will be different (e.g., number of chargers). The proposed approach allows operators to formulate a single query that can be applied to all their charging garages.
@@ 11 Conclusion & Future Work
This work proposes an approach for querying knowledge graphs and signals. While the knowledge graph captures information on the CPS, signals provide a powerful means of modeling run-time sensor data. We have demonstrated the feasibility and applicability of this approach by defining, implementing, and applying the query language SigSPARQL.
The proposed approach enables context-aware monitoring of CPSs, allowing users to pose the same queries across multiple system instances, as knowledge about the concrete system can be accessed within a query. Real-world adoption of such an approach requires further research and development, including a mature and optimized prototype and adaptations of the surrounding technology stack. For example, the SPARQL protocol must be extended to support interactions with continuously running queries and asynchronous updates. These efforts may be justified by the potential benefits in energy efficiency and sustainability that comprehensive monitoring can bring to CPSs.
Concrete possible future directions include (i) designing an optimized prototype for performance evaluations and large-scale monitoring use cases, (ii) adding advanced signal processing functions (e.g., integration), (iii) conducting user evaluations to assess usability benefits of the language, (iv) extending the query language’s semantics to temporal knowledge graphs, and (v) investigating the requirements of real-world CPS deployments and how such an approach can benefit them.